# Differentially Private Matchings

### v2

@@
Differentially Private Matchings
@@
Motivated by the importance of computing matchings in sensitive graph data, we study differentially private algorithms for computing maximum matchings and $b$-matchings in general graphs. Our algorithms return implicit solutions for these problems. We provide a number of algorithms and lower bounds for these problems in different models and settings.
  • We first prove a lower bound showing that computing explicit solutions necessarily incurs large error, even if we try to obtain privacy by allowing ourselves to output non-edges.
  • We then consider implicit solutions, where at the end of the computation there is an ($\varepsilon$-differentially private) billboard and each node can determine its matched edge(s) based on what is written on this publicly visible billboard. For this solution concept, we provide tight upper and lower (bicriteria) bounds, where the degree bound is violated by a logarithmic factor, which we show is necessary via a novel lower bound.
  • We further show that our algorithm can be implemented in the local edge DP (LEDP) model, and can even be done in a logarithmic number of rounds if we further relax the degree bounds by logarithmic factors.
  • Moreover, we implement our algorithms in the node-DP setting by combining edge-DP algorithms with a novel use of arboricity sparsifiers. Interestingly, we prove an impossibility result for publicly releasing such sparsifiers under differential privacy, even though they form a key component of our node-DP algorithm. Our techniques also allow us to improve the bipartite results of [hsu2014private] by a polylogarithmic factor.
@@
  • We demonstrate that all of our results can also be implemented in the continual release model under both edge- and node-privacy.
  • Finally, we extend our main algorithmic result to the $b$-matching setting, including an extension of our node-DP algorithm using a novel approximate $b$-matching arboricity sparsifier. Our sparsifier may be of independent interest even in the non-private graph setting.
1 Introduction
Maximum matching and its variants are central problems within graph algorithms, in both theory and practice. Theoretically, matching and its variants including maximal matching and $b$-matching have been studied in essentially every modern model of computation, including the sequential [edmonds1965maximum, HK71, MV80], streaming [feldman2024maximum, assadi2022semi, kapralov2013better], distributed [CHS09, Fischer20], dynamic [neiman2015simple, BDL21, GP13], and parallel [BFS12, FischerN18] models. In addition, a variety of lower bounds have also been proven [balliu2021lower, konrad2024unconditional].
Matching has been used in practice for many applications on sensitive data including ad allocation [mehta2013online], kidney exchange [BMW22], and online dating [wang2018privacy, yi2016practical]. For each of these applications, the user’s preferences and the existence of an edge between two individuals contain very sensitive information. So it is natural (and important) to study private versions of matching algorithms.
Differential privacy [DMNS06] is the gold standard for protecting the privacy of individuals. While differential privacy was originally introduced in the context of numerical and statistical databases, a growing number of works have placed its focus on sensitive graph data. Since many problems in graph algorithms are actually problems about graph analytics over graphs containing sensitive personal information, there has been a large amount of important work on differentially private graph algorithms. This includes differentially private versions of graph problems like triangle counting, shortest paths, graph cuts, densest subgraph, $k$-core decompositions, and many others [NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13]. In the graph setting, differential privacy is generally defined in terms of edge-neighboring graphs or node-neighboring graphs where the graphs differ in one edge or all edges incident to any one node, respectively; this intuitively corresponds to requiring that when we run our algorithm on neighboring graphs, the distribution of outputs is very similar (see Section 2 for formal definitions).
[…]
Thus, very natural and important questions remain: is it possible to give differentially private algorithms for maximum (or maximal) matchings and $b$-matchings in general graphs beyond bipartite graphs? In this paper, we introduce the study of privately computing matchings and $b$-matchings in general graphs. We give the first algorithms and lower bounds for this question in a variety of models including both edge- and node-DP, local privacy, and continual release via a variety of novel techniques, some of which may be of independent interest in both the privacy and non-private graph algorithms communities. As a corollary of our techniques, we improve the bounds of [hsu2014private] for the node-private bipartite setting.
1.1 Our Contributions
We now state our results and our main technical contributions in the various settings mentioned above.
1.1.1 Solution Types
Explicit Solutions.
It is easy to see that given an input graph, it is impossible to differentially privately output any subgraph (including a matching) containing only true edges; doing so would obviously reveal the existence of any edge that we output. So we must turn to other ways to output a matching that will not result in privacy violations.

### v3

@@
$\AtEveryCite\localrefcontext[sorting=ynt$]
Differentially Private Matchings: Symmetry Lower Bounds, Arboricity Sparsifiers, and Public Vertex Subset Mechanism
Computing matchings in graphs is a foundational algorithmic task. Despite extensive interest in differentially private (DP) graph analysis, work on privately computing matching solutions, rather than just their size, has been sparse. The sole prior work in the standard model of pure $\varepsilon$-differential privacy, by [hsu2014private, STOC’14 ], focused on allocations and was thus restricted to bipartite graphs. This paper presents a comprehensive study of differentially private algorithms for maximum matching and $b$-matching in general graphs, which also yields techniques that directly improve upon prior work in the bipartite setting. En route to solving these matching problems, we develop a set of novel techniques with broad applicability, including a new symmetry argument for DP lower bounds, the first private arboricity-based sparsifiers for node-DP, and the novel Public Vertex Subset Mechanism. We demonstrate the versatility of these tools by applying them to other DP problems, such as vertex cover [GuptaLMRT10, SODA’10], and beyond DP, such as low-sensitivity algorithms [varma2023average, SODA’21,SICOMP’23]. Our contributions include the following:
  • We present the first utility improvements for $\varepsilon$-differentially private bipartite matching since [hsu2014private].
  • We establish a strong lower bound on the error for computing explicit matching solutions, using a novel symmetry argument. This bound demonstrates a fundamental barrier, holding even when the algorithm is permitted to output non-edges. We further generalize this to a lower bound on graph sparsity for synthetic graph generation.
  • Motivated by our lower bound, we introduce a framework for implicit solutions and develop the Public Vertex Subset Mechanism. This new tool allows us to achieve tight (bicriteria) upper and lower bounds for the implicit matching problem.
  • We implement our primary algorithm for implicit matching in the local edge-DP (LEDP) model, achieving logarithmic round complexity.
  • To address the challenging node-DP setting, we introduce the first arboricity-based graph sparsifiers under differential privacy. We demonstrate the independent utility of these sparsifiers by applying them to the node-DP vertex cover and low-sensitivity matching problems.
  • We show that all of our algorithms can be implemented in the continual release model, providing strong guarantees under both edge- and node-privacy.
1 Introduction
Computing matchings is a foundational problem in graph algorithms with broad applications. While decades of research have yielded efficient non-private algorithms [edmonds1965maximum, HK71, MV80, feldman2024maximum, assadi2022semi, kapralov2013better, CHS09, Fischer20, neiman2015simple, BDL21, GP13, BFS12, FischerN18, balliu2021lower, konrad2024unconditional] across numerous computational models, many modern applications, from ad allocation, kidney exchange, to social recommendations, involve sensitive relational data [mehta2013online, BMW22, wang2018privacy, yi2016practical]. For such settings, differential privacy (DP) has emerged as the gold standard for simultaneously providing rigorous privacy guarantees and ensuring good quality solutions. Despite a surge in recent interest in DP graph algorithms and analytics [peng2025synthetic, AHS21, NRS07, GuptaLMRT10, BGM22, DMN23, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, RS16, raskhodnikova2016differentially, Upadhyay13], algorithms that privately compute the matching solution itself, rather than just its scalar value, remain underexplored. The foundational work of [hsu2014private, STOC’14 ] in the central $\varepsilon$-DP model is the sole prior work and was restricted to bipartite graphs for allocation problems. As in the non-private setting, general matching is a harder problem than bipartite matching and requires a new set of techniques.
This paper initiates a comprehensive study of differentially private algorithms for maximum matching and $b$-matching in general graphs; we also give the first utility improvements for the special case of $\varepsilon$-DP bipartite matching since [hsu2014private]. Along the way, we develop broadly useful techniques: a new symmetry argument yielding lower bounds for private explicit solutions, the first arboricity-based sparsifiers under node-DP (together with a novel impossibility proof for their public release), and the Public Vertex Subset Mechanism (PVSM) for privately selecting and announcing differentially private vertex subsets. We show how these tools lead to tight (bicriteria) upper and lower bounds for implicit matching, efficient implementations in the local edge DP (LEDP) model [qin17generating, IMC21locally, IMC21communication, DLRSSY22, ELRS22, MPSL25] with logarithmic communication rounds, and applications to other problems such as vertex cover. We also implement all of our matching algorithms in the continual release model under both edge- and node-privacy, achieving better guarantees compared to previous work.
A maximum matching solution is fundamentally a set of edges. However, releasing any set composed exclusively of edges from the input graph inherently violates differential privacy by directly revealing the existence of those sensitive edges. Therefore, any differentially private algorithm that outputs an explicit matching solution must be permitted to include “non-edges”; that is, given an input graph $G=(V,E)$, it must include edges not present in $E$ in order to ensure that released edges do not just contain those that actually exist. A common relaxation, therefore, is to permit the algorithm to output an explicit matching $H$ on the complete graph $K_{n}$, which may include “non-edges” (edges from $K_{n}\setminus E$). This model is best understood by viewing the input $G$ as a 0/1-weighted complete graph: edges in $E$ have weight 1, and non-edges have weight 0. Edge-DP in this context means neighboring graphs differ by a single edge’s weight (a 0/1 flip). The algorithm outputs an unweighted explicit matching $H$, and its utility is measured by summing the (private) weights of the edges it contains (i.e., its intersection with $E$, $|H\cap E|$). We compare this utility to the non-private maximum matching size, $\operatorname{OPT}$, in $G$. Formally, a $(\gamma,\beta)$-approximation outputs a matching $H$ whose total weight is at least $(\operatorname{OPT}/\gamma)-\beta$.
We prove a fundamental barrier to this approach using a novel symmetry argument. Our lower bound demonstrates two key implications: first, if the output $H$ is constrained to be a matching, it is impossible to achieve any reasonable utility. Second, even if we relax this to allow $H$ to be a $b$-matching (where the degree of every vertex in $H$ can be at most $b$), the degree $b$ must be large to guarantee non-trivial utility. Thus, our lower bound demonstrates a trade-off: to achieve high utility (i.e., include many 1-weight edges), the output $H$ must have a large degree $b$. If $b$ is small (e.g., $b=1$ for a true matching), the utility is necessarily low. The below theorem illustrates a simplified version of our lower bound:
Theorem 1.1 (Simplified Lower Bound for Explicit Solutions; See Theorem 2.1).
Let $\mathcal{A}$ satisfy $(\varepsilon,\delta)$-edge DP, where $\varepsilon=O(1)$ and $\delta=\operatorname{poly}(\nicefrac{{1}}{{n}})$, and always output a graph $H$ of maximum degree at most $b$. If $\mathcal{A}$ returns a $(\gamma,\beta)$-approximation (even in expectation) to the maximum matching on $0/1$-weighted inputs, then:
  • If $b=1$, then either $\gamma=\Omega(n)$ or $\beta=\Omega(n)$, or
  • If $b>1$, then $\beta=\Omega(n-\gamma\cdot b)$ or $\gamma=\Omega\left(\frac{n-\beta}{b}\right)$.
Theorem 1.1 shows that if we restrict $H$ to be a matching, then either the multiplicative or additive error must be $\Omega(n)$. It further shows that even if we allow $H$ to be a $b$-matching for constant $b$, the multiplicative or additive error must still be $\Omega(n)$. Thus, we cannot use the returned graph $H$ as an explicit solution and expect good utility.
Our lower bound thus motivates the study of implicit solutions for matching using the billboard model. The billboard model is a standard framework in differential privacy where a differentially private transcript (the “billboard”) is released publicly. While this public data alone is insufficient to generate a solution to the problem, each node (or “user”) can then privately decode its own high-utility solution (in the case of matching, its matched edges) by combining the public billboard with its own private adjacency list.
We study a natural form of implicit solutions (Definition 3.6) where, roughly speaking, each vertex can implicitly select a subset of its neighbors to be matched with. Even for this particular implicit solution, we show that any $\gamma$-approximate $\varepsilon$-edge DP maximum matching must contain at least one node matched with $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes. This lower bound is also proven using our symmetry argument.
Theorem 1.2 (Simplified Implicit Solution Lower Bound; See Theorem 5.3).
For any $\gamma>1$, a $\gamma$-approximate implicit solution in the billboard model to the maximum matching problem (Definition 3.6), produced by an $\varepsilon$-edge DP algorithm, contains at least one node matched to at least $\Omega\left(\frac{\log(n)}{\varepsilon}\right)$ other nodes, with probability at least $\nicefrac{{1}}{{2}}$.
We emphasize the critical difference between the solution types given in Theorem 1.1 and Theorem 1.2. Theorem 1.1 is a lower bound for explicit solutions in which the publicly released solution itself gives the matching. Theorem 1.2 instead gives a lower bound for implicit solutions that are locally decoded by each node using the public billboard in combination with its own private adjacency list. Hence, Theorem 1.2 is not a trivial consequence of Theorem 1.1, as the set of matched neighbors within an implicit solution is locally decoded by each vertex using the billboard and its adjacency list, rather than explicitly output by the algorithm.
Due to Theorem 1.2, any $\varepsilon$-DP implicit solution with good utility must be a $b$-matching with $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$. There are many settings in which we might want a private matching (or $b$-matching) in a general graph. For example, in online dating, an individual is often provided with a shortlist of top candidates (resulting in a $b$-matching). The matching should be kept private but individuals in every pair should know who they are matched with. We design an algorithm that produces implicit $(2+\eta)$-approximate $b$-matchings for any $b=\Omega\left(\frac{\log(n)}{\varepsilon}\right)$, which is tight against our lower bound for implicit solutions.
Theorem 1.3 (Simplified Implicit $b$-Matching; See Theorem 2.2).
Let $\varepsilon,\eta\in(0,1)$, and $b=\Omega\left(\frac{\log(n)}{\eta^{3}\varepsilon}\right)$. There is an $\varepsilon$-DP algorithm that, with high probability, outputs an implicit $(2+\eta)$-approximate $b$-matching in the billboard model.
This paper establishes a comprehensive set of upper and lower bounds for $\varepsilon$-DP maximum matching and $b$-matching; we present a complete set of novel results across both edge- and node-privacy in the central [DMNS06], local (LEDP) [IMC21locally, DLRSSY22], and continual release models [CSS11, DNPR10]. See Section 1.1 for a summary of our contributions. Our results are driven by three primary novel technical contributions that have broad-reaching applications beyond our paper to other graph algorithms as well.
Symmetry Argument for Graph DP Lower Bounds.
Our lower bound for explicit matchings is driven by a novel symmetry argument inspired by distributed graph algorithms. The core idea is to analyze a “symmetrized” algorithm $\mathcal{A}^{\prime}$ created by wrapping any $(\varepsilon,\delta)$-edge DP algorithm $\mathcal{A}$ as follows: on input $G$, $\mathcal{A}^{\prime}$ computes $\pi^{-1}(\mathcal{A}(\pi(G)))$ for a uniformly random vertex permutation $\pi$. This wrapper preserves $(\varepsilon,\delta)$-DP and worst-case expected utility. We then show that for certain graph problems, such as maximum matching, a large set of random vertex permutations must give equal utility guarantees by definition of DP, leading to high error.
We analyze this on the hard instance family $\mathcal{M}$ of all perfect matchings. When fractional matchings are allowed, symmetrization enforces that, for any $G\in\mathcal{M}$, the expected fractional weight assigned to every non-edge is identical. We then consider a pair of inputs $G_{1},G_{2}\in\mathcal{M}$ which differ by two edges (i.e., given edges $\{u,v\},\{a,b\}\in G_{1}$, create edges $\{u,b\}$ and $\{a,v\}$ in $G_{2}$). The DP guarantee bounds the expected weight on any true edge by $e^{4\varepsilon}$ times the average non-edge weight (plus a $\delta$ term). The degree bound on the returned matching then bounds this average non-edge weight by the algorithm’s degree budget $b$. Summing $\delta$ over the $n/2$ edges of the input matching yields our main lower bound: any $(\gamma,\beta)$-approximation outputting a graph of maximum degree $b$ must satisfy
@@ 1 Introduction
For constant $\varepsilon$ and $\delta\leq 1/n$, this implies a stark trade-off: either $b=\Omega(n)$ or the multiplicative $\gamma$ or additive error $\beta$ must be large (potentially $\Omega(n)$ for constant $b$).
This symmetrization argument extends almost verbatim to the fractional “synthetic graph” relaxation (Definition 4.1), replacing the maximum degree $b$ with the average degree budget $\bar{b}=\frac{1}{n}\sum_{v}b_{v}$:
@@
Thus, any DP algorithm aiming for low error must produce an output synthetic graph with linear average degree. This information-theoretic barrier demonstrates why dense outputs are unavoidable for explicit solutions, explaining the limitations of mechanisms like randomized response and formally motivating our focus on the billboard model with implicit solutions for our upper bounds. We believe our symmetry argument, used to establish multiple lower bounds in our paper, holds significant promise for proving lower bounds to a wider class of DP graph problems beyond matchings.
Public Vertex Subset Mechanism (PVSM).
At the core of our implicit algorithm in the billboard model is the Public Vertex Subset Mechanism (PVSM), a novel technique within graph DP of potential independent interest. Consider an input graph $G=(V,E)$ where each node wants to select a subset of its neighbors as part of the solution to a graph problem. The Public Vertex Subset Mechanism (PVSM) is our new graph DP tool that allows each node to implicitly select a subset of its neighbors. The mechanism’s public output, which is posted to the billboard, does not reveal the specific neighbors chosen. However, it is designed so that each neighbor can combine this public information with their private adjacency list to locally decode whether they were selected, thus ensuring the privacy of the individual edges is maintained.
The mechanism works as follows: first, for each pair of nodes $\{u,v\}$ and parameter $r$, a public coin is flipped with probability $p_{r}=(1+\eta)^{-r}$. These coin flips are all public, collectively defining a set of random subgraphs picked from the complete graph $K_{n}$. A “proposer” node $v$ privately selects an optimal parameter $r^{*}$ (using, e.g., the AboveThreshold mechanism) indicating a subset of neighbors to select. The selection of $r^{*}$ balances utility (capturing many neighbors) with a degree constraint on the output. Crucially, $v$ publicly releases only the index $r^{*}$ to the billboard.
Any “receiver” node $w$ can then locally decode its status using $r^{*}$ and the public coin flips. Node $w$ checks its private adjacency list to see if it is a neighbor of $v$. If it is, $w$ consults the public coin for the pair $\{v,w\}$ at level $r^{*}$. In the context of matchings, if the coin is heads, $w$ knows it has been selected by $v$, thus forming the implicit matched edge. This mechanism provides a general method for privately selecting and publicly encoding graph subsets, and we believe it can be readily adapted to other graph problems in the billboard model. This mechanism yields Theorem 1.3.
Arboricity-Based Node-DP Sparsifiers.
Beyond edge-DP, we also study node-DP, where neighboring input graphs differ by a single node and all its incident edges. We design the first arboricity-based sparsifiers tailored to matching. These sparsifiers (privately) reduce vertex degrees to $\widetilde{O}(\alpha)$ (where $\alpha$ is the arboricity), while maintaining a large enough ($b$-)matching in the sparsified graph. Our sparsifiers guarantee a property known as stability which ensures that the number of edges that differ from any (sparsified) neighboring graph is also bounded by $\widetilde{O}(\alpha)$. The sparsification reduces the edge sensitivity from a worst case $O(n)$ to $\widetilde{O}(\alpha)$. Then, we run our edge-DP algorithms on this sparsified graph to yield high-utility node-DP algorithms for maximum $b$-matching. This yields:
Theorem 1.4 (Simplified Node-DP $b^{\prime}$-Matching; See Theorem 2.6).
Let $\alpha$ denote the arboricity of the input graph $G=(V,E)$. We give a $\varepsilon$-node DP algorithm in the billboard model that, with high probability, outputs an implicit $b$-matching whose size is at least a $(2+\eta)$-approximate $b^{\prime}$-matching, for
$b=O\left(\frac{\alpha\log n}{\eta\varepsilon}+\frac{\log^{2}n}{\eta\varepsilon^{2}}+\frac{b^{\prime}\log n}{\varepsilon}\right).$
Prior work on node-DP sparsification has focused exclusively on reducing graphs using bounded maximum degree (and almost bounded maximum degree)  [BBS13, Day16, kasiviswanathan2013analyzing, wagaman2024time, CZ13]. Our work introduces the first node-DP sparsifier designed around a more general graph property: arboricity. This technical advancement is significant because the class of graphs with bounded arboricity (which includes, for example, all planar graphs and graphs from any minor-closed family) is substantially broader than the class of bounded-maximum-degree graphs.
We demonstrate the independent use of our novel arboricity-based sparsifier by applying it to develop a new node-DP algorithm for vertex cover as well as a new low-sensitivity matching algorithm [varma2023average, yoshida2021sensitivity, yoshida2026low]. We believe arboricity sparsification is of independent interest beyond DP matching.
Improvement on the Utility of [hsu2014private].
For bipartite graphs with supply $s=O\left(\frac{\log^{3.5}(n)}{\varepsilon}\right)$, [hsu2014private] design a node-DP $(1+\eta)$-approximate one-sided $s$-matching algorithm. As an extension of our techniques, we improve the bound to $s=O\left(\frac{\log n}{\varepsilon}\right)$ while maintaining the $(1+\eta)$-approximation guarantee (Theorem 2.8).
1.1 Summary of Results
We give the following set of results which we describe in more detail in Section 2. See Table 1 for a comprehensive summary of quantitative results.
  • Explicit solutions are provably hard. We prove a strong lower bound for computing explicit matching solutions under edge-DP via a new symmetry argument. The bound holds even if the algorithm is allowed to output non-edges and even when we only compare the returned $b$-matching to the size of a (non-private) maximum matching. We further generalize this to a lower bound on the sparsity of differentially private synthetic graph generation.
  • Implicit solutions via PVSM. Motivated by the lower bound, we work in the billboard model that allows for the release of implicit solutions. We introduce the Public Vertex Subset Mechanism (PVSM), a generic mechanism for differentially privately selecting public vertex subsets that are locally decodable by a node’s neighbors using only public coins and their own private adjacency lists. We give tight (bicriteria) upper and lower bounds for implicit maximum matching and $b$-matching in general graphs.
  • Local edge-DP (LEDP) in logarithmic rounds. We implement our implicit matching algorithm in the $\varepsilon$-LEDP model. Our sequential $\varepsilon$-LEDP billboard algorithm achieves optimal dependence on $b=O\left(\frac{\log n}{\varepsilon}\right)$ and a $(2+\eta)$-approximate maximum $b$-matching. We then give a distributed $\varepsilon$-LEDP algorithm that completes in $O(\log n)$ rounds using a proposer/receiver mechanism inspired by non-private distributed algorithms.
  • Node-DP via arboricity sparsification. To handle node-DP, we design the first arboricity-based sparsifiers under differential privacy, and use them to obtain implicit node-DP algorithms for matching in general graphs. We also show an impossibility for the public release of such sparsifiers. As another application of our sparsifier beyond matchings, we obtain a node-DP vertex cover algorithm.
  • Improving [hsu2014private]. We improve the bipartite node-DP guarantees of [hsu2014private] by decreasing $s$ (the one-sided degree parameter of the matching given in [hsu2014private]) from $s=O\left(\frac{\log^{3.5}(n)}{\varepsilon}\right)$ to $s=O\left(\frac{\log(n)}{\varepsilon}\right)$ while maintaining a $(1+\eta)$-approximate $s$-maximum matching.
@@ 1.1 Summary of Results
  • Continual release. All of our results extend to the continual release model for both edge- and node-privacy, with accuracy guarantees that hold for every update.
1.2 Related Work
Private Allocations.
As discussed, the closest work to ours is the private allocation paper by [hsu2014private]. They give a node-differentially private algorithm that implicitly outputs an approximate maximum matching solution on bipartite graphs. The authors model any bipartite graph as a set of bidders (on one side) and goods (on the other). Then, they formulate a differentially private version of the ascending auction price algorithms of [kelso1982job] to compute Walrasian equilibrium prices of the goods (a set of prices ensuring optimal utility of buyers and no good is over-demanded). The algorithm releases differentially private counts of the number of bids each good has received as an implicit solution, using the private counting algorithms of [CSS11, DNPR10]. Assuming the supply $s$ satisfies $s=\Omega(\log^{3.5}(n)/\varepsilon)$, this yields a node-DP $(1+\eta)$-approximate $s$-matching.
Now, auction-based methods and the implicit solution concept of [hsu2014private] are inherently tied to bipartite graphs. While it is possible to reduce general matchings to a bipartite instance via random bipartitions [liu2024pointwise], this transformation incurs a factor of 2 in the multiplicative approximation factor and the obtained approximation ratio is only in expectation. This would imply a node-DP $(2+\eta)$-approximate $s$-matching in expectation for general graphs. Our novel solution concept allows us to handle general graphs, under edge-DP or node-DP assuming low arboricity, and obtain a $(2+\eta)$-approximate $s$-matching with high probability, while reducing the lower bound to $s=\Omega(\log(n)/\varepsilon)$. Moreover, specializing our techniques to the bipartite case (with arbitrary arboricity) recovers the node-DP $(1+\eta)$-approximation ratio of [hsu2014private], while reducing the bound on $s$ to $s=O(\log(n)/\varepsilon)$.
We remark that a followup work [hsu2016jointly] studies a class of private convex programming problems that capture fractional allocations with supply $s$ under approximate $(\varepsilon,\delta)$-differential privacy. They output an $s^{\prime}$-matching with $s^{\prime}>s$, but unlike our results, they compute a fractional allocation and guarantee that the sum of constraint violations ($\sum_{j}s_{j}^{\prime}-s_{j}$) is bounded, whereas our algorithm returns an integral solution while ensuring each constraint violation is bounded ($\max_{j}s_{j}^{\prime}-s_{j}$). Furthermore, our results are given for pure $\varepsilon$-DP (and $\varepsilon$-LEDP). Thus, neither the privacy nor the utility guarantees are comparable to ours.
Private Matching Size.
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [dong2024joins, FHO21, wagaman2024time, raskhodnikova2024fully, epasto2024sublinear] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
Implicit Solutions.
Implicit solutions in the graph setting were first introduced by [GuptaLMRT10] for various combinatorial optimization problems such as set cover and vertex cover. An implicit solution is one where the released information from a differentially private algorithm is $\varepsilon$-DP or $\varepsilon$-LEDP but the released solution does not immediately solve the problem. In particular, the released solution is an implicit solution from which each individual node can decode the explicit solution for itself. In the case of our matching algorithms, the explicit solution computed by every node is the set of edges in the matching that are adjacent to it.
Low-Sensitivity Matching.
Let $\mathcal{A}$ be a randomized graph algorithm that takes as input a graph $G$ and outputs a subset of vertex pairs $F_{G}\subseteq\binom{V}{2}$. Consider edge neighboring graphs $G\sim G^{\prime}$ that differ by a single edge and recall the 1-Wasserstein distance of the output distributions with respect to the symmetric difference metric is given by the following infimum over couplings $\Pi(F_{G},F_{G^{\prime}})$ of the output distributions:
$W_{1}(\mathcal{A}(G),\mathcal{A}(G^{\prime}))=W_{1}(F_{G},F_{G^{\prime}})=\inf_{\pi\in\Pi(F_{G},F_{G^{\prime}})}\left[|(F_{G}\setminus F_{G^{\prime}})\cup(F_{G^{\prime}}\setminus F_{G})|\right]\,.$
[varma2023average] defined the (worst-case) sensitivity of a matching algorithm $\mathcal{A}$ on an input graph $G$ to be
$\max_{e\in E(G)}W_{1}(\mathcal{A}(G),\mathcal{A}(G-e))\,.$
The matching algorithm with the best sensitivity bound for bounded-degree graphs is a $(1+\eta)$-approximation algorithm with sensitivity $\Delta^{O(1/\eta)}$ over graphs of maximum degree $\Delta$ due to [yoshida2026low].
For edge-neighboring graphs $G\sim G^{\prime}$ with arboricity at most $\alpha$, our stable matching sparsifier for bounded-arboricity graphs (Section 7.1.1) maps $G\mapsto H,G^{\prime}\mapsto H^{\prime}$ to edge-neighboring graphs $H\sim H^{\prime}$ with maximum degree $\Delta=O(\nicefrac{{\alpha}}{{\varepsilon}})$ while approximately preserving the maximum matching up a $(1+\eta)$-multiplicative factor. Running the [yoshida2026low] algorithm on the sparsified graph immediately yields a $(1+\eta)$-approximate matching algorithm with sensitivity $\alpha^{\tilde{O}(1/\eta)}$. For graph classes such as planar graphs with constant arboricity but maximum degree $n$, this improves the worst-case sensitivity from $\operatorname{poly}(n)$ to $O(1)$.
2 Technical Overview
We now give a technical overview for the main ideas behind our results and give the formal theorem statements for our upper and lower bounds.
2.1 Lower Bounds for Explicit Solutions (Section 4)
Lower Bounds via Symmetrization (Sections 4.1 and 4.2).
At a very high level, our lower bound for explicit matching solutions (Theorem 2.1) uses an argument similar to the “packing arguments” of [DMN23] for minimum cut. However, instead of using a packing/averaging argument, we use a symmetry argument reminiscent of symmetry arguments in distributed computing (e.g., [kuhn2016local]) and semidefinite programming (e.g., [dam2015semidefinite]). Namely, we argue that any DP algorithm for a collection of input graphs must be highly symmetric. Highly symmetric is defined in terms of the marginal probability of each non-existing edge; we show that the marginal probability of outputting any non-existent edge is identical.